1. the square root of (2-i) is?


2. for a given ,matrix A=[(5,-6),(-3,2)] the matrix P that diagonalizes A is?


3. for a,b is an element of R, let S be a subset of R^2 defined as S={(x,y) is an element of R^3:x+y+axy=b}. Then S is a subspace of R^2 if...?


4. suppose U={(x,y,x+y,z,2y+z) is an element of F^5:x,y,z is an element of F}, then a subspace W of F^5 such that F^5=U denote W is...?


5. the vectors (1,-1,2),(2,3,1),(3,2,t) are not basis of R^3 if...?



1
Expert's answer
2022-01-17T16:54:07-0500

1)


Let "\sqrt{(2-i)}=a+bi" where "a,b\in\R"



"2-i=a^2-b^2+2" abi


"a^2-b^2=2" and "2ab=-1" "\implies\>b=\frac{-1}{2a}"


"a^2-(\frac{-1}{2a})^2=2\implies4a^4-8a^2-1=0"



"a^2=\frac{8^+_-\sqrt{64+16}}{8}" But "a\in\R"


"\therefore\>a^2=2.118"


"a=^+_-1.4553"


"b=\frac{-1}{2a}=^+_-0.3436"


"\sqrt{2-i}=^+_-(1.4553-0.3436i)"



2)


"A=\begin{pmatrix}
 5 & -3 \\
 -6 & 2
\end{pmatrix}"


Characteristic polynomial

"P(t)=det(A-tI)=\begin{pmatrix}
 5-t&-3 \\
 -6&2-t
\end{pmatrix}"


"=(5-t)(2-t)-18"

"=t^2-7t-8"


Solving "t^2-7t-8=0\>,t=8 \>or-1"

Eigenvalues are "\lambda=8\>\>\>\>\>\lambda=-1"


For "(A-8)x=0"

"A-8=\begin{pmatrix}
 -3 & -3 \\
 -6& -6
\end{pmatrix}" "\to^{R_2-2R_1}_{then\>\frac{-1}{3}R_1}" "\begin{pmatrix}
 1 & 1\\
 0& 0
\end{pmatrix}"



"x_1=-x_2"

Eigen vector "=\begin{pmatrix}
 -1 \\
 1 
\end{pmatrix}"


For "(A+1)x=0"

"A+1=\begin{pmatrix}
 6 & -3\\
 -6& 3
\end{pmatrix}\to^{R_2+R_1}_{Then\>\frac{1}{3}R_1}\begin{pmatrix}
 2 & -1 \\
 0 & 0
\end{pmatrix}"



"2x_1=x_2"

Eigenvector = "\begin{pmatrix}
 1 \\
 2 
\end{pmatrix}"


"P=\begin{pmatrix}
 -1 & 1\\
 1 & 2
\end{pmatrix}"









3)

The zero of "\R" is defined as

"0+0+a(0)(0)=0"


This shows that "S" is a subspace of "\R^2" if and only "b=0"


Also "S" is a subspace of "\R^2" if it is closed under addition and scalar multiplication


Let "x_1y_1\in\R^2" and "x_2y_2\in\R^2"

"(x_1+y_1),(x_2+y_2):(x_1+y_1)+(x_2+y_2)+a(x_1+y_1)(x_2+y_2)=0"



"(x_1y_1)+(x_2y_2)=x_1+y_1+ax_1y_1+x_2+y_2+ax_2y_2"


"=(x_1+y)+(y_1+y_2)+a(x_1y_1+x_2y_2)"



"S" is not a subspace of "\R^2" unless "a=0" "b=0"


Let "\alpha\in\R"

"\alpha(x_1y)=\alpha(x+y+axy)"

"=\alpha\>x+xy+a\alpha\>xy"


"\alpha(x)\>x(y)=\alpha\>x+\alpha\>y+a\alpha^2\>xy"


"S" is not a subspace of "\R^2" Unless "a=0,b=0"



4)


"U=x\begin{pmatrix}
 1 \\
 0\\
1\\
0\\
0
\end{pmatrix}+y\begin{pmatrix}
 0 \\
 1 \\
1\\
0\\
2
\end{pmatrix}+z\begin{pmatrix}
 0 \\
 0 \\
0\\
1\\
1
\end{pmatrix}"



"U" is a subspace spanned by linearly independent set "S=\begin{Bmatrix}
( 1,0,1,0,0)+ (0,1,1,0,2)+(0,0,0,1,1)
 
\end{Bmatrix}"


By adjoining elements of "F^5" that will not affect linear independence of

"(0,0,0,1,1)" and "(0,0,0,0,1)"

Span W


"=" { "\begin{bmatrix}
 0 \\
 0 \\
0\\
1\\
1
\end{bmatrix}" ",\begin{bmatrix}
 0 \\
 0 \\
0\\
0\\
1
\end{bmatrix}" }



Or "W=\begin{Bmatrix}
 (0,0,0,x,x+y)\\
 
\end{Bmatrix}"


Such that "F^5" "=u\bigoplus\>w"



5)

The subset "S" of the vector space "V" is not a basis if

(i) S is linearly dependent

(ii) S is not a spanning set



Row echelon form of S is "\begin{pmatrix}
 2&0&0 \\
 0&\frac{7}{2}&0 \\
0&0&\frac{-5t}{7}\frac{+15}{7}
\end{pmatrix}"


"\implies\frac{-5t}{7}+\frac{15}{7}=0"

"\therefore\>t=3"


The vectors are not basis of "\R^3" if "t\ne3"